OpenStudy (anonymous):

Inegrate (x^2)/(x^4 +1)dx

OpenStudy (anonymous):

\[\int\limits_{}^{}x ^{2}dx/\left( x ^{4}+1 \right)\]

OpenStudy (anonymous):

- To integrate this we use quetient rule: =[2x(x^4)-4x^3(x^2)]/(x^4)^2 =[2x^5-4x^5]/x^6 =-2x^5/x^6

OpenStudy (shubhamsrg):

@WONDEMU seems you differentiated it instead of integrating,,that too wrongly! :D

OpenStudy (anonymous):

dude what is that guy doing, I specifically said to integrate

OpenStudy (anonymous):

any of you guys need a hint

OpenStudy (mimi_x3):

Trig sub might work; \(x^2=tan\theta\) => \(2xdx=sec^2\theta\)

OpenStudy (anonymous):

∫ x^2/(x^4 +1) dx = 1/2 ∫ (2x^2) /(x^4+1) dx = 1/2 ∫ (x^2+1 +X^2 -1 ) /(x^4 +1) dx = 1/2 ∫ (x^2 +1 ) /(x^4 + 1) dx + ∫ (x^2 -1 ) /(x^4 +1 ) dx = 1/2 [ A+B] A = ∫ (x^2 +1 ) /(x^4 + 1) dx = ∫ (1 +1/x^2 ) /(x^2 + 1/x^2) dx dividing by x^2 for solving A we can put u=x-(1/x) then we have 1 + 1/x^2 = du also x^2 +1/x^2 = x^2 +1/x^2 -2 +2 = (x -1/x)^2 +2 = u^2 +2 A = ∫ du /(u^2 +2) = 1/√2 arctan ( u /√2) = 1/√2 arctan( (x^2 -1) /(√2x) ) B = ∫ (x^2 -1 ) /(x^4 +1 ) dx ∫ (1 -1/x^2 ) /(x^2 +1/x^2 ) dx for solving B like A we put u=x+(1/x) hence 1 - 1/x^2 = du also x^2 +1/x^2 = X^2+ 1/x^2 +2 -2 = (x +1/x)^2 -2 = u^2 -2 ∫ du /(u^2 -2 ) = 1/2√2 ln ((u - √2) /(u +√2) ) = 1/2√2 ln [ ( x +1/x -√2) / (x+1/x +√2) ] = 1/2√2 ln [ (x^2 -√2 x +1 ) /(x^2 +√2 x+1) ] so the final answer is 1/2 [A+B]+c = 1/2√2 arctan( (x^2 -1) /(√2x) ) + 1/4√2 ln [ (x^2 -√2 x +1 ) /(x^2 +√2 x+1) ]+c c is a constant

OpenStudy (shubhamsrg):

leme try a very indifferent method : split denominator as (x^2 +i)(x^2 -i) where i is sqrt(-1) now multiply divide num by 2 so we have 1/2 [2x^2/(x^2 +i)(x^2 -i) ] dx or 1/2[(x^2 + x^2) /(x^2 +i)(x^2 -i)]dx or 1/2[(x^2 + i + x^2 -i) /(x^2 +i)(x^2 -i)]dx or 1/2[1/(x^2 +i) + 1/(x^2 -i)]dx now integral of 1/(x^2 -a^2) = 1/2a (log(x-a) - log(x+a)) +C..where a is const.. also integral of 1/(x^2 +a^2) = 1/a tan^-1 (x/a) +C same thing goes here.. let m^2 =i so we have : 1/2 [1/2m (log(x-m)/(x+m)) + 1/m (tan^-1 (x/m) ] +C now we just have to simplify x-m/x+m = (x- m)^2 /(x^2 -i) = x^2 +i -2xm /(x^2 -i) = (x^2 +i -2xm)(x^2 +i) /(x^4 -1) = (x^4 + 2x^2 i -2x^3 m -1 -2xim ) /(x^4 -1) also x/m = xm/i = -xmi

OpenStudy (shubhamsrg):

1/m = m/i = -mi so we finally have : 1/2 [-mi/2 log ((x^4 + 2x^2 i -2x^3 m -1 -2xim )/(x^4 -1)) -mi (tan^-1 (-mix) ] +C some knowledge of hyperbolic eqns might be able to bring whole into real terms..hmm

OpenStudy (shubhamsrg):

or even converting to exponential might help..

OpenStudy (mimi_x3):

This might work as well.. \[x^2=\tan\theta => 2xdx=\sec^2d\theta \] \[dx = \frac{\sec^{2}\theta}{2x} d\theta => \frac{\sec^{2}\theta}{2\sqrt{tanx}} d\theta \] \[\int\limits\frac{x^{2}}{x^{4}+1} dx => \int\limits\frac{\tan\theta}{(1+\tan\theta)^{2}} *\frac{\sec^{2}\theta}{2\sqrt{\tan\theta}} \]

OpenStudy (anonymous):

You are somewhat making progress mini_x3

OpenStudy (anonymous):

Although you might want to recheck your last step

OpenStudy (mimi_x3):

whats wrong/

OpenStudy (anonymous):

well if you take x^2 =tan theta will x^4 + 1 be = (1+tan theta)^2 or 1+tan^2 theta

OpenStudy (mimi_x3):

lawl. i made a typo sorry \[\int\limits\frac{x^{2}}{x^{4}+1} dx => \int\limits\frac{\tan\theta}{1+(\tan\theta)^{2}} *\frac{\sec^{2}\theta}{2\sqrt{\tan\theta}} \]

OpenStudy (anonymous):

now at least it might seem a little easier once you turn 1 +tan^2 theta into sec^2 theta

OpenStudy (mimi_x3):

Yes; you can try it. (: After than manipulate it; then two trig subs I think..

OpenStudy (mimi_x3):

that*

OpenStudy (anonymous):

Now here is the greatest problem it comes out to be:\[\int\limits_{}^{}\sqrt{\tan \theta} d \theta \]

OpenStudy (anonymous):

with a 1/2 outside

OpenStudy (mimi_x3):

dont simplify too much!!!

OpenStudy (anonymous):

so basically what I am asking is what is the integral of root over tan theta

OpenStudy (mimi_x3):

First; dont simplify it too much.. convert tanx to sinx/cosx then \(u=sinx-cosx\)

OpenStudy (anonymous):

wait what is the integrand you are getting

OpenStudy (mimi_x3):

@apoorvk: can take over :p

OpenStudy (anonymous):

dude is there anyway to see the document it is very messy